OpenStudy (anonymous):

A right triangle has legs of length 3x m. And 4x m. And a hypotenuse of 75 m. Find the lengths of the legs of the triangle. PLEASE HELP I've got no clue

jimthompson5910 (jim_thompson5910):

a^2 + b^2 = c^2 (3x)^2 + (4x)^2 = 75^2 9x^2 + 16x^2 = 5625 25x^2 = 5625 I'll let you finish

OpenStudy (anonymous):